OpenStudy (anonymous):

Let f(x)=5x+3 and g(x)=2x-5. Find (f+g)(x), (f-g)(x), (fg)(x), and (f/g)(x). Give the domain of each.

OpenStudy (anonymous):

im doing the same thing!

OpenStudy (anonymous):

5x+3+2x-5 =domain 7x-2

OpenStudy (anonymous):

5x+3-(2x-5)

OpenStudy (anonymous):

10x^2-15x+6x-9= 10x^2-9x-9

OpenStudy (anonymous):

f/g= (f/g)(x)= f(x)/g(x) 5x+3/2x-5

OpenStudy (anonymous):

the domains are(\[(\infty,\infty)\]

OpenStudy (anonymous):

did this help